OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). (2 points) Question 1 options: 1) y = alt= x + 7 2) y = -3x + 3 3) y = alt= x + 3 4) y = -3x + 7

OpenStudy (anonymous):

are these the options or do you need to solve for each one?

OpenStudy (anonymous):

The equation in slope intercept form and now need to find the equation parallel to this line and passing through (-1,6). To do this we will use the formula, (y - y1) = m (x - x1) Using this formula with our information we get... (y - 6) = -3 ( x +1) --> distribute the -3 to the x and the 1... y -6= -3x -3 --> y=-3x+3. We now have our equation of the line which is parallel, having the same slope, to the original equation and passes through the point (-1,6).

OpenStudy (anonymous):

Does this make sense?